OpenStudy (anonymous):

Please help! Medal + fan (: 2.The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the force of gravity, and t = time. Solve this equation for g, and use your formula to determine the force of gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 feet. Show all steps in your work.

OpenStudy (dape):

Have you tried to solve for \(g\)?

OpenStudy (anonymous):

I've tried, but I'm not very good at math

OpenStudy (dape):

So this is how to do it, \[d=\frac{1}{2}gt^2\] You want to get everything except g on the left hand side, so multiply both sides by \(2\) to make the \(\frac{1}{2}\) into a 1. \[2d=gt^2\] and then divide by \(t^2\) to cancel it on the right hand side: \[\frac{2d}{t^2}=\frac{gt^2}{t^2}=g\] Now just put in the numbers you were given for time and height (=d=distance in this case)